# Test-negative designs with various reasons for testing: statistical bias and solution

### v2

@@ Abstract
Test-negative designs are widely used for post-market evaluation of vaccine effectiveness, particularly in cases where randomization is not feasible. Differing from classical test-negative designs where only healthcare-seekers with symptoms are included, recent test-negative designs have involved individuals with various reasons for testing, especially in an outbreak setting. While including these data can increase sample size and hence improve precision, concerns have been raised about whether they introduce bias into the current framework of test-negative designs, thereby demanding a formal statistical examination of this modified design. In this article, using statistical derivations, causal graphs, and numerical simulations, we show that the standard odds ratio estimator may be biased if various reasons for testing are not accounted for. To eliminate this bias, we identify three categories of reasons for testing, including symptoms, disease-unrelated reasons, and case contact tracing, and characterize associated statistical properties and estimands. Based on our characterization, we show how to consistently estimate each estimand via stratification. Furthermore, we describe when these estimands correspond to the same vaccine effectiveness parameter, and, when appropriate, propose a stratified estimator that can incorporate multiple reasons for testing and improve precision. The performance of our proposed method is demonstrated through simulation studies.
@@ Introduction
The test-negative design (TND, [1]) refers to a study where patients presenting at healthcare facilities are tested for a disease of interest and categorized as test-positive or test-negative. The intervention effect of interest is then estimated as the odds ratio of intervention comparing test-positive and test-negative patients. As an observational study design, the TND has been increasingly used to study vaccine effectiveness (VE) for infectious diseases such as influenza [2], coronavirus [3], and rotavirus [4], especially when randomized trials are infeasible or unethical. Compared to a cohort design, the TND has the advantage of efficiency and cost-effectiveness, and can potentially correct for bias arising from differential healthcare-seeking behaviors [5].
@@
Despite the emergence of TNDs that include various reasons for testing, appropriate methods to analyze such studies remain an open question. The status quo is to ignore the reasons for testing and perform a standard analysis, e.g., via odds ratios and logistic regression, aggregating the data across all reasons for testing. However, this method can be biased for estimating the overall VE since different reasons for testing may involve heterogeneous populations, across which the VE, the likelihood of testing, and covariate distributions can all differ. [14] pointed out that reasons other than symptoms for testing may introduce bias, and [15] conceptualized this issue through a causal graph. Both papers and [16] stated that a stratified analysis based on testing reasons could alleviate the bias, but no complete quantification or formal solution was given. [13] addressed the bias from a naive TND analysis when there were two reasons for testing: (i) symptomatic self-reporting or (ii) testing close contacts of known cases, and proposed a hybrid design to remove such bias. However, they made a strong assumption that the VE for either recruitment population is the same; in addition, other testing reasons, such as routine screening, were not considered in their setting. Finally, a naive solution is to filter out asymptomatic subjects and only use the remaining data to perform a classical TND analysis. Although this procedure is valid, it loses a large amount of information.
@@
In this paper, we combine directed acyclic graphs (DAGs, [17]) with additional assumptions to examine the statistical properties of TNDs with various reasons for testing, and propose several estimands with corresponding unbiased estimators. First, we revisit the classical TND to clarify underlying assumptions and an appropriate estimand, and then identify two additional categories of reasons for testing: disease-unrelated reasons and contact tracing, for which we construct corresponding DAGs and derive the target estimands. The target estimands for these three categories are generally distinct, which expand the classical scope of TNDs and provide new perspectives on the generalizability of TND estimates. Second, given multiple reasons for testing, we explain why the odds ratio estimator pooling all testing reasons may be biased. Finally, we propose several estimators that can target each VE of interest using the corresponding data. Furthermore, we provide conditions to justify a combined estimator based on multiple reasons for testing, thereby maximizing the use of relevant data to achieve unbiased estimation, valid hypothesis testing, and increased precision. Moreover, our proposed methods do not solely depend on parametric model assumptions; they also offer the flexibility to incorporate non-parametric kernel smoothing to approximate the underlying data-generating distribution. To the best of our knowledge, we provide the first vehicle to clarify to what extent non-symptomatic reasons for testing can be included in TNDs without introducing bias, and our proposed methods are the first to accommodate a range of reasons for testing, which can inform statistical practice for future TNDs.
@@
We consider TNDs to study a specific disease, namely Covid-19, with a view to estimating vaccine effectiveness. For simplicity, we consider assessment of only a single vaccination dose, although the development can be extended to cover situations where $V$ captures a more general vaccination history. Thus, for each subject, let $V$ be the vaccination status with $V=1$ if vaccinated and 00 otherwise. Let $I$ be the true infection status for severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2), where $I=1$ if the subject is infected and $I=0$ otherwise, as determined by a definitive diagnostic test. Let $R$ be a categorical variable encoding the reasons for testing with mutually exclusive categories, with each category associated with a single reason, such as “routine screening for work”, or a pool of related reasons, such as “Covid-19-like symptoms”. For a traditional test-negative design, we let $T$ be a binary indicator of being tested at a healthcare setting (and thus eligible for inclusion in a standard test-negative study) with $T=1$ if tested and 00 otherwise. Throughout, we assume all the above variables are measured without misclassification. Finally, $\boldsymbol{X}$ denotes observed covariates, such as age, sex, and calendar time, and $H$ is a latent (unobserved) healthcare-seeking behavior indicator. Again, for simplicity, we assume $H$ is a binary random variable with $H=1$ indicating healthcare-seeking behavior for a symptomatic participant and $H=0$ otherwise. This setup also applies to studies on other infectious diseases, such as influenza, and other non-vaccine interventions with the only difference being that the reasons for testing may be different.
In a standard TND, all eligible patients with Covid-19-like symptoms who are tested are included by design and subsequently classified by whether they test positive for the disease in question—here infection with SARS-CoV-2—or not. This test is assumed to have perfect sensitivity and specificity. This enrollment scheme results in an observational study [1, 18]; the sampling scheme can be considered as a historical cohort study (as vaccination has occurred in the past) or as a form of case-cohort sampling.
A primary goal is inference on the conditional VE for strata encoded by observed covariates, which is simply one minus the conditional relative risk of infection given observed covariates,
$VE(\boldsymbol{x})=1-\dfrac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x})}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x})}.$ (1)
Of note, the conditional VE is a function of $\boldsymbol{x}$, which means inference is made on each stratum, such as a specific age group. VE across different strata can be linked by assuming, for example, that $\Pr(I=1|V,\boldsymbol{X})$ follows a log-binomial regression model
$\log\{\Pr(I=1|V,\boldsymbol{X})\}=\beta_{0}+\beta_{V}V+\beta_{\boldsymbol{X}}^{\top}\boldsymbol{X},$
where $\beta_{0},\beta_{V},\beta_{\boldsymbol{X}}$ are regression coefficients. Under this model, $VE(\boldsymbol{x})=1-e^{\beta_{V}}$ does not vary by $\boldsymbol{x}$, which can be interpreted as reflecting homogeneous VE across strata.
At first glance, $H$ is of less obvious interest for conditional VE estimation. However, $H$ may be associated with both $V$ and $I$, and in fact, TND studies were originally conceived as a way of eliminating, or at least reducing, this potential confounding bias by focusing only on the stratum with $H=1$. Thus, another estimand of interest is the conditional VE for strata defined by observed covariates $\boldsymbol{X}$ and the latent healthcare-seeking behavior $H$, defined as
$VE(\boldsymbol{x},h)=1-\dfrac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},H=h)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},H=h)}.$ (2)
Compared to $VE(\boldsymbol{x})$ defined in Equation (1), $VE(\boldsymbol{x},h)$ further depends on the latent variable $H$ and is hence more challenging to estimate. As noted, the standard TND study automatically stratifies on $H$ as no data is observed for those with $H=0$. The two estimands $VE(\boldsymbol{x})$ and $VE(\boldsymbol{x},h)$ are connected through the collapsibility of relative risks [19]. Mathematically, if we assume healthcare-seeking behavior is not an effect modifier, i.e., $VE(\boldsymbol{x},1)=VE(\boldsymbol{x},0)$, and $V$ is independent of $H$ given $\boldsymbol{X}$, then $VE(\boldsymbol{x})=VE(\boldsymbol{x},h)$.
The two VE estimands characterize associations instead of causality. We next consider their links to causal estimands under structural causal assumptions.
Connection to casual interpretation
We define $I(v)$ as the potential outcome of a subject if their vaccination status were $V=v$ for $v=0,1$. For a causal interpretation of $VE(\boldsymbol{x},h)$, we define
[…]
To provide a causal interpretation for $VE(\boldsymbol{x})$, we have two approaches. The first approach assumes that healthcare-seeking is not a confounder to obtain
${VE(\boldsymbol{x})=\textup{causal-}VE(\boldsymbol{x})=1-\dfrac{\Pr(I(1)=1\mid\boldsymbol{X}=\boldsymbol{x})}{\Pr(I(0)=1\mid\boldsymbol{X}=\boldsymbol{x})}}$
under the above causal assumptions for $VE(\boldsymbol{x},h)$. A second approach assumes the same causal assumptions for $VE(\boldsymbol{x},h)$ and uses proximal learning [21] to address the issue of unmeasured confounders by leveraging measured proxy variables such as vaccine history.
A concise review of statistical analyses for classical TNDs
To illustrate the underlying causal structure of TNDs, we use directed acyclic graphs (DAGs, [17]), which describe the relationships (arrows) among variables (nodes) and are widely used in medical and epidemiological research. Of note, the DAG itself does not encode all necessary assumptions for TNDs [22], and we combine DAGs and additional assumptions for our derivation (See Table 1). Figure 1 shows a DAG for classical TNDs where infection-related symptoms are the only reason for testing. Here, we additionally denote $S$ as an indicator of having Covid-19-like symptoms, with $I=1$ and $S=1$ together implying being diseased. By definition, $R=\textup{Symptoms}$ implies $S=1$. We assume that patients cannot self-diagnose based on symptoms, which is reflected in Figure 1 by no $I\rightarrow R=\textup{Symptoms}$ or $I\rightarrow T$ direct effects. In addition, an important feature of classical TNDs is that enrolled individuals are all healthcare-seekers who complete the test and have Covid-19-like symptoms, i.e., those with $H=1$, $S=1$, and $T=1$, thereby providing no information on VE among healthcare non-seekers or untested individuals.
$\boldsymbol{X}$$V$$I$$S=1$ $R=$ Symptoms $T=1$$H=1$
@@
Beyond the DAG, we assume that vaccination does not reduce non-SARS-CoV-2 causes of symptoms among healthcare-seekers [1], i.e.,
${\Pr(I=0,S=1|V=1,\boldsymbol{X},H=1)=\Pr(I=0,S=1|V=0,\boldsymbol{X},H=1).}$
These assumptions together motivate an odds-ratio estimator for the VE on the relative risk scale:
@@
$={1-\frac{\Pr(I=1,S=1|V=1,\boldsymbol{X},H=1)}{\Pr(I=1,S=1|V=0,\boldsymbol{X},H=1)}.}$ (3)
@@
We denote the right-hand side of Equation (3) as $VE_{s}(\boldsymbol{X},1)$, as it evaluates VE for symptomatic infections among healthcare-seekers. Here, standard TNDs achieve VE identification via odds ratio because of its unique design that controls are also healthcare seekers with Covid-like symptoms and that vaccination has no effect on other types of infections, which together cancels out the collider bias from $T$.
In addition, $VE_{s}(\boldsymbol{X},1)$ may differ from $VE(\boldsymbol{X},1)$ since $VE(\boldsymbol{X},1)$ also accounts for vaccination effects on the risk of asymptomatic infections. This relationship can be made explicit mathematically. Specifically, if $RR_{P}$ is the relative risk of progression to the defined symptom criteria conditional on infection, comparing vaccinated to unvaccinated individuals, then
${VE_{s}(\boldsymbol{X},1)=1-[1-VE(\boldsymbol{X},1)]RR_{P}}.$
The difference between $VE(\boldsymbol{X},1)$ and $VE_{s}(\boldsymbol{X},1)$ thus depends on $RR_{P}$, which in turn is potentially influenced by the definition of symptoms required for recruitment.
To require equivalence of $VE_{s}(\boldsymbol{X},1)$ and $VE(\boldsymbol{X},1)$, we thus need an additional restrictive assumption that vaccination has no effect on preventing symptom development given infection, i.e., $RR_{P}=1$ or, equivalently,
${\Pr(S=1|V=1,I=1,\boldsymbol{X},H=1)=\Pr(S=1|V=0,I=1,\boldsymbol{X},H=1).}$
Although this assumption is known to be violated for Covid-19 vaccines, it may be plausible if $V$ represents non-vaccine interventions, e.g., wearing facial masks [6], social distancing, or other hygiene behaviors. Given this assumption, classical TNDs can target $VE(\boldsymbol{x},1)$. However, vaccination often leads to symptom alleviation so that $VE_{s}(\boldsymbol{x},1)$ is usually larger than $VE(\boldsymbol{x},1)$, indicating increased effectiveness for preventing symptomatic infections.
Two additional categories of reasons for testing
We next consider two common types of reasons for testing other than symptoms and discuss their target VE. We name the two categories of reasons as “disease-unrelated reasons” and “case contact tracing”, each of which is described with a DAG as presented in Figures 2(a) and 2(b), respectively. For these two categories, we assume a cohort sampling scheme, which better reflects the actual recruitment procedure than case-cohort sampling and facilitates the statistical identification.
$\boldsymbol{X}$$H$$V$$I$$S$ $R=$ Disease- unrelated reasons $T=1$
$\boldsymbol{X}$$H$$V$$I$$S$ $R=$ Case contact tracing $T=1$
@@
Figure 2: DAGs for testing due to disease-unrelated reasons (a) and case contact tracing (b). Circle nodes stand for random variables, while rectangle nodes represent that the random variables are conditioned on specific values.

### v3

@@ Abstract
Test-negative designs are widely used for post-market evaluation of vaccine effectiveness, particularly in cases when randomized trials not feasible. Differing from classical test-negative designs where only healthcare-seekers with symptoms are included, recent test-negative designs have involved individuals with various reasons for testing, especially in an outbreak setting. While including these data can increase sample size and hence improve precision, concerns have been raised about whether they introduce bias into the current framework of test-negative designs, thereby demanding a formal statistical examination of this modified design. In this article, using statistical derivations, causal graphs, and numerical demonstrations, we show that the standard odds ratio estimator may be biased if various reasons for testing are not accounted for. To eliminate this bias, we identify three categories of reasons for testing, including symptoms, mandatory screening, and case contact tracing, and characterize associated statistical properties and estimands. Based on our characterization, we show how to consistently estimate each estimand via stratification. Furthermore, we describe when these estimands correspond to the same vaccine effectiveness parameter, and, when appropriate, propose a stratified estimator that can incorporate multiple reasons for testing and improve precision. The performance of our proposed method is demonstrated through simulation studies.
@@ Introduction
The test-negative design (TND, [1]) refers to a study where patients presenting at healthcare facilities are tested for a disease of interest and categorized as test-positive or test-negative. The intervention effect of interest is then estimated as the odds ratio of intervention comparing test-positive and test-negative patients. As an observational study design, the TND has been increasingly used to study vaccine effectiveness (VE) for infectious diseases such as influenza [2], coronavirus [3], and rotavirus [4], especially when randomized trials are infeasible or unethical. Compared to a cohort design, the TND has the advantage of efficiency and cost-effectiveness, and can potentially reduce bias arising from differential healthcare-seeking behaviors [5].
@@
Despite the emergence of TNDs that include various reasons for testing, appropriate methods to analyze such studies remain an open question. The status quo is to ignore the reasons for testing and perform a standard analysis, e.g., via odds ratios and logistic regression, aggregating the data across all reasons for testing. However, this method can be biased for estimating the overall VE since different reasons for testing may involve heterogeneous populations, across which the VE, the likelihood of testing, and covariate distributions can all differ. [14] pointed out that reasons other than symptoms for testing may introduce bias, and [15, 16] conceptualized this issue through causal graphs. Both papers and [17] stated that a stratified analysis based on testing reasons could alleviate the bias, but no complete quantification or formal solution was given. [13] addressed the bias from a naive TND analysis when there were two reasons for testing: (i) symptomatic self-reporting or (ii) testing close contacts of known cases, and proposed a hybrid design to remove such bias. However, they made a strong assumption that the VE for either recruitment population is the same; in addition, other testing reasons, such as mandatory screening, were not considered in their setting. Finally, a naive solution is to filter out asymptomatic subjects and only use the remaining data to perform a classical TND analysis. Although this procedure is valid, it loses a large amount of information.
@@
In this paper, we combine directed acyclic graphs (DAGs, [18]) with additional assumptions to examine the statistical properties of TNDs with various reasons for testing, and propose several estimands with corresponding unbiased estimators. First, we revisit the classical TND to clarify underlying assumptions and an appropriate estimand, and then identify two additional categories of reasons for testing: mandatory screening and contact tracing, for which we construct corresponding DAGs and derive the target estimands. The target estimands for these three categories are generally distinct, which expand the classical scope of TNDs and provide new perspectives on the generalizability of TND estimates. Second, given multiple reasons for testing, we explain why the odds ratio estimator pooling all testing reasons may be biased. Finally, we propose several estimators to target each VE of interest using the corresponding data. Furthermore, we provide conditions to justify a combined estimator based on multiple reasons for testing, thereby maximizing the use of relevant data to achieve unbiased estimation, valid hypothesis testing, and increased precision. Moreover, our proposed methods do not solely depend on parametric model assumptions; they also offer the flexibility to incorporate non-parametric kernel smoothing to approximate the underlying data-generating distribution. To our knowledge, we provide the first vehicle to clarify to what extent asymptomatic reasons for testing can be included in TNDs without introducing bias, and our proposed methods are the first to accommodate a range of reasons for testing, which can inform statistical practice for future TNDs.
@@
We consider TNDs to study a specific disease, namely Covid-19, with a view to estimating vaccine effectiveness. For simplicity, we consider assessing only a single vaccination dose, although the development can be extended to cover situations where $V$ captures a more general vaccination history. Thus, for each subject, let $V$ be the vaccination status with $V=1$ if vaccinated and 00 otherwise. Let $I$ be the true infection status for severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2), where $I=1$ if the subject is infected and $I=0$ otherwise, as determined by a definitive diagnostic test. Let $S$ be an indicator of having Covid-like symptoms, such as fever and coughing. We use $\boldsymbol{X}$ to denote observed confounders, such as age, sex, and calendar time, and $H$ is a latent (unobserved) healthcare-seeking behavior indicator. Again, for simplicity, we assume $H$ is a binary random variable with $H=1$ indicating actively seeking care if feeling ill. Finally, we define $C$ as a binary indicator of having contact with known cases, which can be impacted by the vaccination status.
For TND studies involving multiple reasons for testing, we define $R$ as a categorical variable encoding the reasons for testing with mutually exclusive categories, with each category associated with a single reason, such as “routine screening for work”, or a pool of related reasons, such as “Covid-like symptoms”. We let $T$ be a binary indicator of being tested at a healthcare setting (and thus eligible for inclusion in a standard TND) with $T=1$ if tested and 00 otherwise. Throughout, we assume all the above variables are measured without misclassification. This setup also applies to studies on other infectious diseases, such as influenza, and other non-vaccine interventions, such as enforcing social distance, with the only difference being that the reasons for testing may be different.
We assume the joint distribution of the defined random variables is characterized by the DAG in Figure 1. For the pre-test variables $\mathcal{Z}=\{X,H,V,C,I,S\}$, the assumed causal structure is the same as TND literature [5, 15, 16, 19] with the only difference being the new post-vaccination variable $C$, which can directly impact infection and symptoms. We introduce this variable for situations when testing is due to case contact tracing. For testing-related variables $(R,T)$, since $R$ may involve multiple categories of reasons, the joint distribution of $(R,T)$ can depend on all pre-test variables as shown in Figure 1. However, for a specific reason $r$, $P(R=r,T=1|\mathcal{Z})$ may be relevant to only a subset of $\mathcal{Z}$, and we characterize three categories in the next section.
$X$$V$$I$$S$$H$$C$$R$$T$Pre-test variables $\mathcal{Z}=\{X,H,V,C,I,S\}$
Figure 1: The data generating process described by a DAG. The blue arrow is a simplified presentation of arrows from each node in $\mathcal{Z}$, e.g., $\mathcal{Z}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\rightarrow}R$ stands for six arrows from all pre-test variables to $R$. Here, $X,H,V,C,I,S,R,T$ are observed confounders, healthcare-seeking behavior, vaccination, case contact, infection, symptom, reason for testing, and testing, respectively.
In a TND involving multiple reasons for testing, all individuals tested for SARS-CoV-2 are included in the study. In a prospective design, individuals are recruited, tested, and then categorized as cases or controls. Alternatively, a retrospective design examines electronic health records and extracts data related to testing for analysis. For both implementations, we regard the individual data $(\mathcal{Z},R)$ with $T=1$ as random samples from the joint distribution characterized by Figure 1. Furthermore, our framework also accommodates a meta-analysis scheme, where individual data are first sampled from $\Pr(\mathcal{Z}|R=r,T=1)$ for each reason $r$ and then data across testing reasons are pooled for analysis.
Three categories of reasons for testing
Testing due to symptoms as in classical TNDs
The classical TND only involves individuals who seek care and are tested for infection because of Covid-like symptoms, and we denote them with $R=\textup{Symptoms}$. Therefore, we assume that
$\Pr(R=\textup{Symptoms},T=1|\mathcal{Z})=H\times S\times f_{s}(\boldsymbol{X},V)$
for some unknown function $f_{s}$, which means testing due to symptoms implies $H=1$ and $S=1$ (by definition) and can further depend on observed confounders and vaccination status. Here, we assume testing is not driven by the contact history or the true infection status, which implies that this conditional probability is independent of $C$ or $I$. For example, $f_{s}(\boldsymbol{X},V)=0.3-0.1V$ means 30% chance of testing due to symptoms for unvaccinated symptomatic healthcare seekers, but this probability is reduced by 10% for vaccinated individuals. With this assumption on $(R=\textup{Symptoms},T=1)$, the DAG in Figure 1 implies the DAG for testing due to symptoms in Figure 2.
$X$$V$$I$$S=1$$H=1$$C$$R=\textup{Symptoms},T=1$
@@
Beyond the DAG, we assume that vaccination does not reduce non-SARS-CoV-2 causes of symptoms among healthcare-seekers, i.e.,
${\Pr(I=0,S=1|V=1,\boldsymbol{X}=\boldsymbol{x},H=1)=\Pr(I=0,S=1|V=0,\boldsymbol{X}=\boldsymbol{x},H=1).}$
This assumption is commonly used in standard TNDs to reduce an odds ratio parameter to relative risk [1, 19]. These assumptions together motivate an odds-ratio estimator for the VE on the relative risk scale:
@@
$={1-\frac{\Pr(I=1,S=1|V=1,\boldsymbol{X}=\boldsymbol{x},H=1)}{\Pr(I=1,S=1|V=0,\boldsymbol{X}=\boldsymbol{x},H=1)}.}$ (1)
@@
We denote the right-hand side of Equation (1) as $VE_{symp}(\boldsymbol{x},1)$, as it evaluates the conditional VE against symptomatic infections, i.e. being diseased, among healthcare-seekers. Of note, the conditional VE is a function of $\boldsymbol{x}$, which means inference is made on each stratum, such as a specific age group. Here, classical TNDs achieve VE identification via odds ratio because of its unique design that controls are also healthcare seekers with Covid-like symptoms and that vaccination has no effect on other types of infections, which together cancels out the collider bias from $T$ and target the relative risk.
Mandatory screening
Mandatory screening refers to testing due to massive screening, back-to-school prerequisites, international traveling, or other factors that are irrelevant to their infection status or symptoms. In other words, tests are motivated by external requirements instead of internal concerns about health. By this definition, there is no direct effect from $H,C,I,S$ to $R=\textup{Mandatory screening}$ or $T=1$, which implies
${\Pr(R=\textup{Mandatory screening},T=1|\mathcal{Z})=f_{m}(\boldsymbol{X},V)}$
for some unknown function $f_{m}$. Here, mandatory screening can be directly impacted by observed covariates and vaccination status since unvaccinated subjects may be subject to stricter testing requirements.
Given this assumption and Figure 1, we obtain the DAG for mandatory screening in Figure 3 and
$1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Mandatory screening},T=1)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Mandatory screening},T=1)}$
$=1-\dfrac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x})}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x})},$ (2)
where the left-hand side is an estimable quantity from data under mandatory testing, and the right-hand side defines the conditional VE against infections (covering both symptomatic or asymptomatic infections), which we denote as $VE(\boldsymbol{x})$.
$X$$V$$I$$S$$H$$C$$R=\textup{Mandotory screening},T=1$
@@
Figure 3: The DAG for testing due to mandatory screening. Circle nodes stand for random variables, while rectangle nodes represent that the random variables are conditioned on specific values.
If Covid-like symptoms are also measured in the data from mandatory screening, following a similar procedure, we can also estimate the conditional VE against symptomatic infections by
$1-\frac{\Pr(I=1,S=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Mandatory screening},T=1)}{\Pr(I=1,S=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Mandatory screening},T=1)}$
$=1-\dfrac{\Pr(I=1,S=1|V=1,\boldsymbol{X}=\boldsymbol{x})}{\Pr(I=1,S=1|V=0,\boldsymbol{X}=\boldsymbol{x})}.$ (3)
We denote the VE at the right-hand side of Equation (3) as $VE_{symp}(\boldsymbol{x})$, which differs for $VE_{symp}(\boldsymbol{x},1)$ in classical TNDs by covering healthcare non-seekers.
Case contact tracing
Case contact tracing is a common practice of monitoring the spread of diseases in an outbreak setting by testing all contacts of known confirmed cases. Because of the mandatory nature of such contact tracing, testing is unrelated to healthcare-seeking behavior, infection status, or symptoms, but it implies $C=1$ and can be related to covariates and vaccination status. Therefore, we assume
$\Pr(R=\textup{Case contact tracing},T=1|\mathcal{Z})=C\times f_{c}(\boldsymbol{X},V)$
for some unknown function $f_{c}$. Given these assumptions, we obtain the DAG for case contact tracing in Figure 4 and have
$1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},C=1,R=\textrm{Case contact tracing},T=1)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},C=1,R=\textrm{Case contact tracing},T=1)}$
$=1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},C=1)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},C=1)},$ (4)
where the left-hand side refers to an estimable quantity from data and the right-hand side represents the conditional VE against infections among contacts of known cases, which we denote as $VE_{cct}(\boldsymbol{x})$.
$X$$V$$I$$S$$H$$C=1$$R=\textup{Case contact tracing},T=1$
Figure 4: The DAG for testing due to case contact tracing. Circle nodes stand for random variables, while rectangle nodes represent that the random variables are conditioned on specific values.
The above procedure can also be adapted to target the conditional VE against symptomatic infections among contacts of known cases by
$1-\frac{\Pr(I=1,S=1|V=1,\boldsymbol{X}=\boldsymbol{x},C=1,R=\textrm{Case contact tracing},T=1)}{\Pr(I=1,S=1|V=0,\boldsymbol{X}=\boldsymbol{x},C=1,R=\textrm{Case contact tracing},T=1)}$
$=1-\frac{\Pr(I=1,S=1|V=1,\boldsymbol{X}=\boldsymbol{x},C=1)}{\Pr(I=1,S=1|V=0,\boldsymbol{X}=\boldsymbol{x},C=1)},$ (5)
which we denote as $VE_{symp,cct}(\boldsymbol{x})$.
Summary
Overall, Table 1 summarizes the three reasons for testing, their assumptions, and the target VE discussed above. Under minimum assumptions, each category of reasons for testing targets its own VE estimand. In addition, classical TNDs are constrained to healthcare seekers, whereas “mandatory screening” and “case contact tracing” cover both healthcare-seekers and non-seekers. Categorization of reasons for testing thus demonstrates the bias from an aggregate odds ratio estimator and also provides a stepping stone for strategically combining reasons for testing, which we introduce below.
In the Web Appendix B, we provide the assumptions needed for a causal interpretation of each VE estimand. Among them, it is worth noting that $VE_{cct}(\boldsymbol{x})$ conditions on a post-randomization variable $C$, and hence can be interpreted as the controlled direct effect in causal mediation analysis [20], instead of the standard causal VE as in $VE(\boldsymbol{x})$.
Table 1: A summary of considered reasons for testing with their assumptions, examples, and target VEs in TNDs. The VE parameters $VE_{symp}(\boldsymbol{x},1)$, $VE(\boldsymbol{x})$, $VE_{symp}(\boldsymbol{x})$, $VE_{cct}(\boldsymbol{x})$, $VE_{symp,cct}(\boldsymbol{x})$ and $VE_{cct}(\boldsymbol{x})$ are defined in Equations (1), (2), (3), (4), and (5), respectively.
Reasons for testing Main assumptions Examples Target VEs
Symptoms (Figure 2) 1. Only symptomatic healthcare-seekers      are tested. 2. Vaccination does not reduce     test-negative symptoms. 3. Given a symptom, the uptake of testing     is independent of true infection status     or contact with cases. Fever, fatigue VE against symptomatic infections for healthcare seekers: $VE_{symp}(\boldsymbol{x},1)$.
Mandatory screening (Figure 3) 1. Testing is not directly impacted     by infection status, symptom, or     healthcare-seeking behavior. Massive tests, traveling VE against infection: $VE(\boldsymbol{x})$, or VE against symptomatic infection: $VE_{symp}(\boldsymbol{x})$.
Case contact tracing (Figure 4) 1. Only case contacts are tested. 2. Testing is not directly impacted by     infection status, symptom, or     healthcare-seeking behavior. Case contact tracing VE against infection among case contacts: $VE(\boldsymbol{x})$, or VE against symptomatic infection among case contacts: $VE_{symp}(\boldsymbol{x})$.
Beyond the three testing reasons we discussed, $R$ may encode more categories, for which meaningful VE identification may be challenging. For mathematical completeness on $\Pr(R,T|\mathcal{Z})$, we define $R=\textup{Others}$ as the combination of all other testing reasons and denote $\Pr(R=\textup{Others},T=1|\mathcal{Z})=f_{o}(\mathcal{Z)}$ for some unknown function $f_{o}$, which imposes no structure on testing due to other reasons. Finally, we require $f_{s},f_{m},f_{c},f_{o}$ are non-negative functions with summations uniformly upper-bounded by 1 so that $\Pr(R,T|\mathcal{Z})$ is well defined.
Potential bias given various reasons for testing
When a TND involves multiple reasons for testing (including the above-discussed three categories or other reasons), the DAG in Figure 1 can not be simplified. Since the observed data only contain tested individuals, a naive odds ratio estimator with observed data will be biased for VE, e.g.,